OpenStudy (anonymous):

Simplify the expression if not possible write simplified 1. 2(11d-5) 2. 4(8n+10d-7d)

OpenStudy (anonymous):

10d-7d = how many d's?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

10d - 7d = 3d

OpenStudy (anonymous):

wat howd u get tht

OpenStudy (anonymous):

2. 4(8n+3d) then you could use foil unless they want you to leave it factored...

OpenStudy (anonymous):

10 - 7 = 3

OpenStudy (anonymous):

10 - 8 = 2

OpenStudy (anonymous):

10d's - 7d's = 3d's

OpenStudy (amistre64):

is this a system of equations?

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

no, it's not a system...

OpenStudy (amistre64):

"4(8n+3d); then you could use foil unless they want you to leave it factored..." what is foil?

OpenStudy (anonymous):

i dont even kno wat a system of equation is

OpenStudy (amistre64):

a system of equation is where you have multiple equations and want to find out the points they have in common

OpenStudy (anonymous):

FOIL is a method for multiplying two binomials.. it seems many people will say foil when they mean distribute

OpenStudy (anonymous):

ohhh well thts not wat it is

OpenStudy (amistre64):

:) its not really a method either; just a memory device. might as well use FLOI or OFLI

OpenStudy (anonymous):

idk wat tht is either

OpenStudy (anonymous):

If I understand your question correctly: It looks like you can sove for d in the first equation, and substitue that into the 2nd equation. \[2(11d-5)=0\]\[22d-10=0\]\[22d=10\]\[d=10/22\]\[d=5/11\]Now plug into 2nd equation\[4[n+10(5/11)-7(5/11)]=0\]\[4[n+(50/11)-(35/11)]=0\]\[4n+(200/11)-(140/11)=0\]\[4n=(140/11)-(200/11)\]\[4n=-60/11\]\[n=-60/44\]\[n=-15/11\]

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I would say it's both ;).. but yes FOIL is a mnemonic

OpenStudy (amistre64):

by posting 2 questions at once; with same variables; its easy to confuse us as towards the nature of the problem :)

OpenStudy (anonymous):

I'm pretty sure they wanted those simplified individually

OpenStudy (anonymous):

ya i did